a)how many cubic inches in 273 cubic ft?
b)how many cubic yards in 273 cubic feet?

Answer by checkley77(12844)

1 cubic foot=12*12*12=1,728 cubic in.
273*1,728=471,744 cubic in. in 273 cubic ft.
1 cubic yard=3*3*3=27 cubic ft.
273/27=10.111 cubic yd. in 273 cubic ft.